Water rises in a capillary tube to a certain height such that the upward force due to surface tension is balanced by 75 × 10-4 N, forces due to the weight of the liquid. If the surface tension of water is 6 × 10-2 N/m, the inner-circumference of the capillary must be
(a) 1.25 × 10-2 m (b) 0.50 × 10-2 m
(c) 6.5 × 10-2 m (d) 12.5 × 10-2 m

Grade:12

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option d

Regards
Vikas TU
3 years ago
Upwards force due to surface tension,
Fu=(2πR)×T
where, 2πR→circumference of tube
Fu=(2πR)×T,
Because surface tension is force per unit length,
⇒75×10^−4=(2πR)×6×10^−2
(2πR)=75×10^−4 / 6×10^−2=12.5×10^−2m